# Iterated INLA for State and Parameter Estimation in Nonlinear Dynamical Systems

### v1

@@ 1 Introduction
Physics-based modelling plays a major role in science and engineering even in today’s landscape of machine learning, where heavy emphasis is placed on data-driven modelling. In applications, such as numerical weather prediction (NWP), the number of observations received daily, while plentiful, pales in comparison to the sheer dimensionality of the state—typically of the order of $\mathcal{O}(10^{9})$, while the number of observations is of the order $\mathcal{O}(10^{7})$ Office [2024]. Under such data-scarce regimes, it is crucial to incorporate expert knowledge into models so that we can extrapolate in regions outside of the data distribution while equipping them with sound uncertainty estimates (Figure 1).
@@
Figure 1: Comparison of the predictions made from a non-physics-informed model (GPR) vs. a physics-informed model (iterated INLA). Gray dots are observations.
@@
Alternatively, there has been a recent surge of interest in integrating physical knowledge into machine learning models. Physics-informed neural networks (PINNs) introduced by Raissi et al. [2019] for example, use neural networks to parametrically represent the model state to estimate the state and parameters given data, by minimising a tailored loss function comprising a data fit term and a PDE residual term. Analogous ML-based methods using Gaussian processes (GPs) have also been proposed, such as AutoIP [Long et al., 2022], which replaces the neural network in PINNs by GPs to equip predictions with uncertainties using variational inference. A similar method has been proposed by Chen et al. [2021] that also provides convergence guarantees in the limit of increasing collocation points. However, all these methods encode the PDE knowledge artificially through the likelihood, instead of directly embedding them in the prior, which makes interpretability of their predicted uncertainties harder. On the other hand, latent force models [Alvarez et al., 2009], or the recent work by Nikitin et al. [2022] injects ODE/PDE knowledge directly into the GP prior by designing physics-informed kernels. However, their approach is limited to modelling linear PDEs, such as the heat and wave equations.
We overcome the limitations posed above by proposing a DA method inspired by SPDE-INLA [Rue et al., 2013]. This is a statistical inference technique that represents GP priors by a Gaussian Markov Random Field (GMRF), enabling the posterior marginals on the state and model hyperparameters to be inferred efficiently using Integrated Nested Laplace Approximation (INLA) [Rue et al., 2009]. We extend their method to handle nonlinear PDE priors by iteratively linearising the PDE to produce GP approximations to the prior at each iteration, and subsequently using INLA to update the state and parameter estimates. In particular, this allows us to (1) use priors derived from arbitrary nonlinear PDEs without having to encode them in the likelihood, and (2) provide accurate non-Gaussian estimates to the state and posterior marginals, going beyond the setting of Gaussian approximations, necessarily imposed in variational inference or ensemble Kalman methods.
@@ 2 Background
where (1) expresses the differential equation satisfied by $u(t,\cdot)$ in the interior of a compact domain $D$, perturbed by a spatio-temporal noise term $\dot{\mathcal{W}}$; (2) expresses the boundary conditions, and (3) describes the initial condition, which we model as a Gaussian random element with mean $u_{b}$ and covariance $\mathcal{C}$. Denoting by $V$ a topological vector space where the solution $u(t,\cdot)$ to (1)–(3) lives for $t\in[0,T]$, we take $\mathcal{M}$ in (1) to be a possibly nonlinear operator on $V$ that may additionally depend on a set of parameters $\boldsymbol{\theta}$ that are a priori unknown. Equations (1)–(3) encode the prior knowledge we have about our state $u(t,x)$.
@@ 2.1 Ensemble and Variational DA
In general, the distributions (7)–(9) can be approximated arbitrarily well using sequential Monte-Carlo (SMC) techniques [Chopin et al., 2020]. However, this can be expensive to compute and moreover suffers from a so-called “weight-collapse" when $d_{u}>\!\!>1$, making it unreliable to use in high-dimensional settings [Bengtsson et al., 2008]. The ensemble Kalman filter/smoother (EnKF/S) [Evensen, 1994, Evensen and Van Leeuwen, 2000] has been proposed as an appealing alternative in high dimensions, which, like SMC, uses particles to empirically approximate the state distribution, but differs from it by only using information about the first two moments of the particles to condition on data. This effectively employs a Gaussian approximation to all distributions arising in the computations, making it more akin to the extended Kalman filter/smoother (ExKF/S). However, by taking the number of particles to be much smaller than $d_{u}$, computations in EnKF/S can be performed much more efficiently than in ExKF/S, enabling its use in high-dimensional problems such as weather forecasting. However, using a small number of particles can result in inaccurate uncertainty estimates [Bannister, 2017]. Moreover, when we jointly infer the parameters via state-augmentation [Evensen, 2009], parameters of the process noise cannot be inferred accurately [DelSole and Yang, 2010].
@@ 3.2 Nonlinear setting
4:     $\mathcal{L}^{(n)}_{0}\leftarrow$ Linearise operator $\mathcal{L}(\boldsymbol{\theta})$ around $u^{(n)}$
@@
14:  $\boldsymbol{\mu}^{(\infty)}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta}),\boldsymbol{v}_{\boldsymbol{u}|\boldsymbol{y}}^{(\infty)}(\boldsymbol{\theta})$
@@ 3.2.1 Known parameters
Below, we show that updating $\boldsymbol{u}_{0}^{(n)}$ according to (30) is a sound choice, as it is identical to using the Gauß–Newton method to optimise the weak-constraint 4D-Var cost (10).
@@
The damped update of the linearisation point $\boldsymbol{u}_{0}^{(n)}$ in (30) is equivalent to minimising the weak-constraint 4D-Var cost (10) using Gauß–Newton.
@@
This implies firstly, that we have guaranteed convergence of our algorithm in the same setting where the Gauß–Newton method converges (e.g. Theorem 10.1 in Nocedal and Wright [1999]), and secondly, we can interpret the output of Algorithm 1 as the marginals of an approximate Laplace approximation to $p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})$ that is close to the true Laplace approximation when the model is weakly nonlinear (see Appendix B.3).
@@ 3.2.2 Unknown parameters
Updating the linearisation point $\boldsymbol{u}^{(n)}_{0}$ according to (36)–(39) is an approximate Gauß–Newton method for minimising the parameter-averaged 4D-Var cost $\mathbb{E}_{p(\boldsymbol{\theta}|\boldsymbol{y})}[-\log p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})]$.
@@
4:     $\mathcal{L}^{(n)}_{0}\leftarrow$ Linearise operator $\mathcal{L}(\boldsymbol{\theta})$ around $u^{(n)}_{0}$
@@
19:  $\left\{\tilde{p}^{(\infty)}(u_{i}|\boldsymbol{y})\right\}_{i=1}^{d_{u}},\left\{\tilde{p}^{(\infty)}(\boldsymbol{\theta}|\boldsymbol{y})\right\}_{j=1}^{|\boldsymbol{\theta}|}$
@@ 4.1 Stochastic nonlinear pendulum
As baselines, we considered vanilla RBF-GP regression, the ensemble Kalman smoother (EnKS) and AutoIP [Long et al., 2022]. For EnKS, we use the state-augmentation approach [Evensen, 2009] to jointly infer the state and model parameters $(b,c,\sigma_{u})$. However, this method does not accommodate learning of the observation noise $\sigma_{y}$, so we fix this to the ground truth value in the EnKS experiments. AutoIP is capable of learning all four parameters, however it can only learn point estimates by gradient descent. Therefore, we initialise them with fixed values, set to the mode of the respective priors. For GPR, we only learn the hyperparameters of the RBF kernel by type-II maximum likelihood estimation [Rasmussen and Williams, 2006].
@@
In Figure 2, we plot the state uncertainties produced by SMC, AutoIP, EnKS and iINLA-II on a single random seed. Here, we can see that the uncertainties generated by iINLA-II is nearly identical to the SMC output. This is also reflected in the lower MMD score. We also display the estimates of the parameters in Figure 3, where we plot a heatmap of the estimated distribution on the parameters $b$ and $\sigma_{u}$, computed using (a) iINLA-II, and (b) EnKS. As a reference, we also display the marginals on the parameters $b,\sigma_{u}$ computed using SMC in blue. We see that both methods achieve similar results to SMC for estimating $b$. However, for $\sigma_{u}$, we see that while the estimates from iINLA-II agree closely with the results from SMC, the estimate from EnKS is significantly different. This behaviour is consistent with previous observations that ensemble methods struggle to learn parameters associated with stochastic terms in the equation [DelSole and Yang, 2010]. Iterated INLA in contrast can get accurate estimates on the stochastic parameters. As a reference, to produce the SMC results in Figure 3 took $\approx 25$ minutes on an M1 Macbook Pro, whereas it took $\approx 90$ seconds to produce analogous results using iINLA.
@@ 4.2 PDE benchmarks
EnKS $\mathbf{0.008\pm 0.001}$ $\mathbf{-3.67\pm 0.11}$ $\mathbf{0.028\pm 0.001}$ $\mathbf{-4.08\pm 0.076}$ $0.228\pm 0.029$ $-0.010\pm 0.263$
@@
iINLA-I $\mathbf{0.009\pm 0.001}$ $-3.49\pm 0.39$ $0.053\pm 0.003$ $-2.30\pm 0.60$ $\mathbf{0.010\pm 0.000}$ $\mathbf{-3.28\pm 0.03}$
iINLA-II $\mathbf{0.009\pm 0.001}$ $-3.49\pm 0.34$ $0.053\pm 0.004$ $-2.95\pm 0.14$ $\mathbf{0.010\pm 0.000}$ $\mathbf{-3.28\pm 0.04}$
@@
We compared the performance of iINLA against the same baselines of GPR, EnKS and AutoIP, and their results are summarised in Table 2. Generally, we find that iINLA and EnKS perform better than the other two models on both metrics. AutoIP tends to produce over-smoothed results and fails to learn the correct parameter, leading to an MNLL that is even worse than GPR’s. The differences between type I and II updates in iINLA were negligible here.
For the Burgers’ experiment, the performance of EnKS and iINLA are similar, with the EnKS slightly ourperforming on the MNLL. However, we encountered numerical stability issues with the EnKS using a fourth-order Runge–Kutta scheme with a timestep of $\Delta t=0.02$ when jointly learning the state and parameters (we did not encounter this issue when learning just the state). Hence, this required us to use a more sophisticated solver in Kassam and Trefethen [2005] with an order of magnitude smaller timestep of $\Delta t=10^{-3}$ to run the simulations reliably. Iterated INLA did not have this issue and ran reliably at the original timestep, using a basic central difference scheme for discretisation. On the Allen-Cahn example, we see that EnKS outperforms iINLA on both metrics. To understand this, it helps to see that the uncertainties generated by iINLA is generally higher than those generated by EnKF (Figure 4). This is due to the existence of the small but positive process noise $\sigma_{u}$ that cannot be removed from iINLA. Upon training, this converged to $10^{-3}$. Hence, even in regions where predictions can be more confident, the uncertainty cannot go below this value, leading to slighly smoother and underconfident predictions (note that the uncertainty in EnKS goes down to $10^{-5}$). In the KdV example, we instead see that iINLA performs better than EnKS. Again, we encountered numerical stability issues with EnKS on the KdV example and is likely the cause for the poor performance of EnKS. On the other hand, we found that iINLA is numerically robust and converges consistently, without the need for grid upsampling. We plot the outputs of all methods for each PDE in Appendix E.
@@
Figure 4: Comparison of the predicted standard deviations on the Allen-Cahn example. The predictions are generally underconfident for iINLA due to the presence of $\sigma_{u}>0$.

### v2

# Code Availability

@@ 1 Introduction
Physics-based modelling plays a major role in science and engineering even in today’s landscape of machine learning, where heavy emphasis is placed on data-driven modelling. In applications, such as numerical weather prediction (NWP), the number of observations received daily, while plentiful, pales in comparison to the sheer dimensionality of the state—typically of the order of $\mathcal{O}(10^{9})$, while the number of observations is of the order $\mathcal{O}(10^{7})$ Office [2024]. Under such data-scarce regimes, it is crucial to incorporate expert knowledge into models so that we can extrapolate in regions outside of the data distribution (Figure 1), while equipping them with sound uncertainty estimates.
@@
Figure 1: Comparison of the predictions made from a non-physics-informed model (Gaussian process regression / GPR) vs. a physics-informed model (iterated INLA). The ground truth is a simulation of the 1D Burgers’ equation. Gray dots are observation locations.
@@
Alternatively, there has been a recent surge of interest in integrating physical knowledge into machine learning (ML) models. Physics-informed neural networks (PINNs) introduced by Raissi et al. [2019] for example, use neural networks to parametrically represent the model state to estimate the state and parameters given data, by minimising a tailored loss function comprising a data fit term and a PDE residual term. Analogous ML-based methods using Gaussian processes (GPs) have also been proposed, such as AutoIP [Long et al., 2022], which replaces the neural network in PINNs by GPs to equip predictions with uncertainties using variational inference. A similar method has been proposed by Chen et al. [2021] that also provides convergence guarantees in the limit of increasing collocation points. However, all these methods encode the PDE knowledge artificially through the likelihood, instead of directly embedding them in the prior, which makes interpretability of their predicted uncertainties harder. On the other hand, latent force models [Alvarez et al., 2009], or the recent work by Nikitin et al. [2022] injects ODE/PDE knowledge directly into the GP prior by designing physics-informed kernels. However, their approach is limited to modelling linear PDEs, such as the heat and wave equations.
We overcome the limitations posed above by proposing a DA method inspired by the work Rue et al. [2013]. This is a statistical inference technique that represents GP priors by a Gaussian Markov Random Field (GMRF), enabling the posterior marginals on the state and model hyperparameters to be inferred efficiently using Integrated Nested Laplace Approximation (INLA) [Rue et al., 2009]. We extend their method to handle nonlinear PDE priors by iteratively linearising the PDE to produce GP approximations to the prior at each iteration, and subsequently using INLA to update the state and parameter estimates. In particular, this allows us to (1) use priors derived from arbitrary nonlinear PDEs without having to encode them in the likelihood, and (2) provide accurate non-Gaussian estimates to the state and posterior marginals, going beyond the setting of Gaussian approximations, necessarily imposed in variational inference or ensemble Kalman methods.
@@ 2 Background
where (1) expresses the differential equation satisfied by $u(t,\cdot)$ in the interior of a compact domain $D$, perturbed by a spatio-temporal noise term $\dot{\mathcal{W}}$; (2) expresses the boundary conditions for a given operator $\mathcal{B}$ (e.g., the identity map or spatial derivatives), and (3) describes the initial condition, which we model as a Gaussian random element with mean $u_{b}$ and covariance $\mathcal{C}$. Denoting by $V$ a topological vector space where the solution $u(t,\cdot)$ to (1)–(3) lives for $t\in[0,T]$, we take $\mathcal{M}$ in (1) to be a possibly nonlinear operator on $V$ that may additionally depend on a set of parameters $\boldsymbol{\theta}$ that are a priori unknown. Equations (1)–(3) encode the prior knowledge we have about our state $u(t,x)$.
@@ 2.1 Ensemble and Variational DA
In general, the distributions (7)–(9) can be approximated arbitrarily well using sequential Monte-Carlo (SMC) techniques [Chopin and Papaspiliopoulos, 2020]. However, this can be expensive to compute and moreover suffers from a so-called “weight-collapse" when $d_{u}>\!\!>1$, making it unreliable to use in high-dimensional settings [Bengtsson et al., 2008]. The ensemble Kalman filter/smoother (EnKF/S) [Evensen, 1994, Evensen and Van Leeuwen, 2000] has been proposed as an appealing alternative in high dimensions, which, like SMC, uses particles to empirically approximate the state distribution, but differs from it by only using information about the first two moments of the particles to condition on data. This effectively employs a Gaussian approximation to all distributions arising in the computations, making it more akin to the extended Kalman filter/smoother (ExKF/S). However, by taking the number of particles to be much smaller than $d_{u}$, computations in EnKF/S can be performed much more efficiently than in ExKF/S, enabling its use in high-dimensional problems such as weather forecasting. However, using a small number of particles can result in inaccurate uncertainty estimates [Bannister, 2017]. Moreover, when we jointly infer the parameters via state-augmentation [Evensen, 2009], parameters of the process noise cannot be inferred accurately [DelSole and Yang, 2010].
@@ 3.2 Nonlinear setting
4:     $\mathcal{L}^{(n)}_{0}\leftarrow$ Linearise operator $\mathcal{L}$ around $u^{(n)}$
@@
15:  return  $\boldsymbol{\mu}^{(\infty)}_{\boldsymbol{u}|\boldsymbol{y}}(\boldsymbol{\theta}),\boldsymbol{v}_{\boldsymbol{u}|\boldsymbol{y}}^{(\infty)}(\boldsymbol{\theta})$
@@ 3.2.1 Known parameters
Below, we show that updating $\boldsymbol{u}_{0}^{(n)}$ according to (30) is a sound choice, as it is identical to using the Gauss–Newton method to optimise the weak-constraint 4D-Var cost (10).
@@
The damped update of the linearisation point $\boldsymbol{u}_{0}^{(n)}$ in (30) is equivalent to minimising the weak-constraint 4D-Var cost (10) using Gauss–Newton.
@@
This implies firstly, that we have guaranteed convergence of our algorithm in the same setting where the Gauss–Newton method converges (e.g. Theorem 10.1 in Nocedal and Wright [1999]), and secondly, we can interpret the output of Algorithm 1 as the marginals of an approximate Laplace approximation to $p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})$ that is close to the true Laplace approximation when the model is weakly nonlinear (see Appendix B.3).
@@ 3.2.2 Unknown parameters
Updating the linearisation point $\boldsymbol{u}^{(n)}_{0}$ according to (36)–(39) is an approximate Gauss–Newton method for minimising the parameter-averaged 4D-Var cost $\mathbb{E}_{p(\boldsymbol{\theta}|\boldsymbol{y})}[-\log p(\boldsymbol{u}|\boldsymbol{y},\boldsymbol{\theta})]$.
@@
4:     $\mathcal{L}^{(n)}_{0}\leftarrow$ Linearise operator $\mathcal{L}$ around $u^{(n)}_{0}$
@@
20:  return  $\left\{\tilde{p}^{(\infty)}(u_{i}|\boldsymbol{y})\right\}_{i=1}^{d_{u}},\left\{\tilde{p}^{(\infty)}(\boldsymbol{\theta}|\boldsymbol{y})\right\}_{j=1}^{|\boldsymbol{\theta}|}$
@@
Remark 3.4.
The computational cost of iterated INLA is $\mathcal{O}(N_{i}I)$, where $N_{i}$ is the number of iterations and $I$ is the complexity of one interation of INLA (see Section A.1 for more details). There are no significant differences in the costs between type I and II updates. In general, this is cheaper than running particle MC, which requires a large number of particles to accurately estimate the state and parameter posteriors. However, it is more costly than running EnKS, which only scales linearly in the number of time steps and cubically in the ensemble size – the latter is typically chosen to be small.
@@ 4.1 Stochastic nonlinear pendulum
As baselines, we considered vanilla RBF-GP regression (GPR), the ensemble Kalman smoother (EnKS) and AutoIP [Long et al., 2022]. For EnKS, we use the state-augmentation approach [Evensen, 2009] to jointly infer the state and model parameters $(b,c,\sigma_{u})$. We also consider an iterative extension of EnKS (iEnKS) proposed in Bocquet and Sakov [2013], which can be used for joint state and paramter estimation. However, these methods do not accommodate learning of the observation noise $\sigma_{y}$, so we fix this to the ground truth value in the EnKS / iEnKS experiments. AutoIP is capable of learning all four parameters, however it can only learn point estimates by gradient descent. Therefore, we initialise them with fixed values, set to the mode of the respective priors. For GPR, we only learn the hyperparameters of the RBF kernel by type-II maximum likelihood estimation [Rasmussen and Williams, 2006].
@@
In Figure 2, we plot the state uncertainties produced by SMC, AutoIP, EnKS and iINLA-II on a single random seed. Here, we can see that the uncertainties generated by iINLA-II is nearly identical to the SMC output. This is also reflected in the lower MMD score. We also display the estimates of the parameters in Figure 3, where we plot a heatmap of the estimated distribution on the parameters $b$ and $\sigma_{u}$, computed using (a) iINLA-II, and (b) EnKS. As a reference, we also display the marginals on the parameters $b,\sigma_{u}$ computed using SMC in blue. We see that both methods achieve similar results to SMC for estimating $b$. However, for $\sigma_{u}$, we see that while the estimates from iINLA-II agree closely with the results from SMC, the estimate from EnKS is significantly different. This behaviour is consistent with previous observations that ensemble methods struggle to learn parameters associated with stochastic terms in the equation [DelSole and Yang, 2010]. Iterated INLA in contrast can get accurate estimates on the stochastic parameters. As a reference, to produce the SMC results in Figure 3 took $\approx 25$ minutes on an M1 Macbook Pro, whereas it took $\approx 1$ minute to produce analogous results using iINLA (Table 2).
@@
Method Run time (s)
SMC (10,000 samples) $1541\pm 38$
iINLA-II (25 iterations) $67.26\pm 0.94$
EnKS (100 ensembles) $3.07\pm 0.31$
Table 2: Comparison of run times between SMC, iterated INLA (type II) and EnKS to produce the parameter estimates in Figure 3. We display the mean and standard deviation of run times across five different runs on an M1 Macbook Pro.
@@ 4.2 PDE benchmarks
EnKS $0.008\pm 0.001$ $-3.67\pm 0.11$ $\mathbf{0.028\pm 0.001}$ $\mathbf{-4.08\pm 0.076}$ $0.228\pm 0.029$ $-0.010\pm 0.263$
iEnKS $\mathbf{0.006\pm 0.001}$ $\mathbf{-3.97\pm 0.05}$ $0.062\pm 0.002$ $11.67\pm 1.18$ $0.131\pm 0.021$ $2807\pm 650$
@@
iINLA-I $0.009\pm 0.001$ $-3.49\pm 0.39$ $0.053\pm 0.003$ $-2.30\pm 0.60$ $\mathbf{0.010\pm 0.000}$ $\mathbf{-3.28\pm 0.03}$
iINLA-II $0.009\pm 0.001$ $-3.49\pm 0.34$ $0.053\pm 0.004$ $-2.95\pm 0.14$ $\mathbf{0.010\pm 0.000}$ $\mathbf{-3.28\pm 0.04}$
@@
We compared the performance of iINLA against the same baselines of GPR, EnKS, iEnKS and AutoIP. Their results are summarised in Table 3. Generally, we find that iINLA and EnKS perform better than the other models on both metrics (with the exception of the Burgers’ experiment, where iEnKS performs marginally better). AutoIP tends to produce over-smoothed results and fails to learn the correct parameter, leading to an MNLL that is even worse than GPR’s. The differences between type I and II updates in iINLA were negligible here.
For the Burgers’ experiment, the performance of EnKS, iEnKS and iINLA are similar, with the iEnKS slightly ourperforming the others on both the RMSE and the MNLL. However, we encountered numerical stability issues with the EnKS and iEnKS using a fourth-order Runge–Kutta scheme with a timestep of $\Delta t=0.02$ when jointly learning the state and parameters (we did not encounter this issue when learning just the state). Hence, this required us to use a more sophisticated solver in Kassam and Trefethen [2005] with an order of magnitude smaller timestep of $\Delta t=10^{-3}$ to run the simulations reliably. Iterated INLA did not have this issue and ran reliably at the original timestep, using a basic central difference scheme for discretisation. On the Allen-Cahn example, we see that EnKS outperforms iINLA on both metrics (iEnKS performed significantly worse on this example). To understand this, it helps to see that the uncertainties generated by iINLA is generally higher than those generated by EnKF (Figure 4). This is due to the existence of the small but positive process noise $\sigma_{u}$ that cannot be removed from iINLA. Upon training, this converged to $10^{-3}$. Hence, even in regions where predictions can be more confident, the uncertainty cannot go below this value, leading to slighly smoother and underconfident predictions (note that the uncertainty in EnKS goes down to $10^{-5}$). In the KdV example, we instead see that iINLA performs better than EnKS. Again, we encountered numerical stability issues with EnKS on the KdV example and is likely the cause for the poor performance of EnKS. On the other hand, we found that iINLA is numerically robust and converges consistently, without the need for grid upsampling. We plot the outputs of all methods for each PDE in Appendix E.
@@
Figure 4: Comparison of the predicted standard deviations on the Allen-Cahn example. The predictions are generally underconfident for iINLA due to the presence of $\sigma_{u}>0$. Gray dots are observation locations.